OpenStudy (anonymous):

Atmospheric pressure P decreases as altitude h increases. At a temperature of 15°C, the pressure is 101.3 kilopascals (kPa) at sea level, 87.1 kPa at h = 1 km, and 74.9 kPa at h = 2km. Use a linear approximation to estimate the atmospheric pressure at an altitude of 3 km. ___________ kPa

OpenStudy (anonymous):

101.3/87.1 = 14.2 87.1/74.9 = 12.2 each gap was 1km (14.2+12.2)/2 = average pressure decrease per kilometer (linearly) = 13.2 101.3 - (13.2 * 3) = 61.7kPa